OpenStudy (anonymous):

A canoe rental company charges $12 per hour for a canoe. Each person in the canoe is required to wear a life vest. The company charges a flat rate of $3 per life vest. If two people rented a canoe for three hours, how much did they pay? A. $30 B. $33 C. $39 D. $42

OpenStudy (221emily):

D.

OpenStudy (221emily):

12 times 3 =36+ 3 times 2+ 42 do it in a calculator

OpenStudy (anonymous):

thanks

OpenStudy (221emily):

12 times 3 =36+ 3 times 2= 42 do it in a calculator

OpenStudy (221emily):

I messed up the + is =